OpenStudy (anonymous):

Simplify: (3y − 4)^2

OpenStudy (anonymous):

FOIL

OpenStudy (vishweshshrimali5):

(a-b)^2 = a^2 + b^2 - 2ab put a = 3y b = 4

OpenStudy (anonymous):

\[(a - b)^2 = a^2 + b^2 - 2ab\] Here a = 3y and b= 4, plug in the values and then solve..